A satellite is in orbit around Earth at a height of 120 km above Earth’s surface. Find the orbital speed of the satellite. (Mass of the earth is 6*1024 kg and the radius of Earth is 6371 km)
Answer #1

Solution)

Given,

Height, h=120 km

Mass of Earth, Me=6*10^24 kg

Radius, R=6371 km

Here,

Mv^2/(R+h)= GMem/(R+h)^2

So, v^2= GMe/(R+h)

V^2=(6.67*10^-11)*(6*10^24)/(6371+120)*1000

Velocity, v=7900 m/s

============